### v1

@@ Abstract
Large Language Models (LLMs) are increasingly used in software development to generate functions, such as attack detectors, that implement security requirements. However, LLMs struggle to generate accurate code, resulting, e.g., in attack detectors that miss well-known attacks when used in practice. This is most likely due to the LLM lacking knowledge about some existing attacks and to the generated code being not evaluated in real usage scenarios. We propose a novel approach integrating Retrieval Augmented Generation (RAG) and Self-Ranking into the LLM pipeline. RAG enhances the robustness of the output by incorporating external knowledge sources, while the Self-Ranking technique, inspired to the concept of Self-Consistency, generates multiple reasoning paths and creates ranks to select the most robust detector. Our extensive empirical study targets code generated by LLMs to detect two prevalent injection attacks in web security: Cross-Site Scripting (XSS) and SQL injection (SQLi). Results show a significant improvement in detection performance compared to baselines, with an increase of up to 71%pt (percentage points [^1] A percentage points is the standard unit of measure for differences between percentages.) and 37%pt in the F2-Score for XSS and SQLi detection, respectively.
Keywords:
Large Language Models Code Security Attack Detection Retrieval Augmented Generation
@@ 1 Introduction
To this end, we adopt two novel components to the LLM pipeline: Retrieval Augmented Generation (RAG) lewis2020retrieval and Self-Ranking. RAG, a technique widely investigated in Natural Language Processing (NLP), enhances the quality of the output by incorporating external information. We use RAG to leverage existing knowledge sources that document known attacks to increase the robustness of the detectors generated by LLMs. Additionally, LLMs may exhibit non-deterministic behaviour in generation ouyang2023llm , which results in multiple, diverse solutions for the same prompt. We propose Self-Ranking, building upon the concept of Self-Consistency wang2023selfconsistency : we take advantage of the multiple reasoning paths associated with LLM’s non-determinism to rank the alternative outputs and select the best one. More concretely, the LLM can propose different implementations obtained querying the model multiple times, with Self-Ranking we propose a strategy to rank the different implementations and to keep only the best-performing ones. In turn, ranking is based on the creation of a synthetic dataset, also generated by LLMs (hence the moniker Self-Ranking), used to automatically assess the robustness of each candidate output.
@@
In our empirical study, we target two well-known and prevalent attacks: XSS and SQLi HYDARA2015170 ; lawal2016systematic , which are ranked second and third in ‘2023 CWE Top 25 Most Dangerous Software Weaknesses’ [^3] https://cwe.mitre.org/top25/archive/2023/2023_top25_list.html respectively. Our extensive empirical study involves the evaluation of nine different LLMs. Results indicate that integrating external knowledge with RAG improved detection performance up to 66%pt (on average 17%pt) on the F2-Score for XSS detection and up to 67%pt (on average 7%pt) for SQLi detection compared to relying solely on few-shot examples. Additionally, employing Self-Ranking enhanced the LLM performance by up to 71%pt (on average 37%pt) on the F2-Score for XSS detection and up to 43%pt (on average 6%pt) for SQLi detection.
@@
State-of-the-art (SOTA) machine-learning based attack detectors require a labeled training set of attacks, while our LLM-based approach is applicable even when no training set is available, provided a RAG source is accessible. In our empirical study, we considered also the scenario when a training set is available and SOTA methods can be used. In such a scenario, our approach achieved a performance comparable to SOTA XSS and SQLi machine learning-based detection methods, with the remarkable advantage that LLMs are pre-trained, while specialized machine learning-based methods require dedicated model design and training: developers can obtain SOTA performance by just querying an LLM using our RAG and Self-Ranking augmented pipeline. Another key advantage is that LLMs generate code that can be understood and manipulated by developers, while the decision making logic of a machine learning model is to a large extent opaque to developers. We also demonstrated the transferability capabilities of LLMs, which are structurally missing in the SOTA detectors: after configuring the LLM (i.e., setting parameters such as the model checkpoint to use, the temperature, the number of few-shot examples) for optimal performance on one attack detection task (e.g., XSS), the resulting configuration can be transferred to another task (e.g., SQLi), with 16%pt and 10% improvement when transferring the LLM parameters from one task to another, as compared to average cases without parameter transfer.
@@
  • We introduce a novel approach that integrates RAG and Self-Ranking to evaluate and improve the robustness of LLM-generated attack detectors.
@@ 3.2 Few-shot Examples
$>$detect_xss(’http://anonymised-site/moderate-social-media&t=1396533765893&n=1129109&k=mainentity’)
@@
$>$detect_xss(’http://anonymised-site/guestbook/index.php?lang="><script>alert(document.cookie)</script>...’)
@@ 3.3 RAG
We integrate RAG to enhance the LLM’s ability to generate robust detectors. We followed the general RAG methodology, starting from the indexing phase. Once the knowledge sources about attacks are stored in vector representation [^5] The selection of appropriate data sources and their storage is discussed in Section 4 as it is an experimental choice., they can be queried to retrieve the most relevant information chunks (referred to as ‘External Knowledge’ in Figure 1). We utilize the Task (‘Detector Signature’ in Figure 1) directly as a query to retrieve these chunks, which are then appended to the Template to obtain a RAG Prompt, as illustrated in Listing 5.
@@ 3.4 Self-Ranking
Self-Consistency does not apply directly to our task as proposed originally wang2023selfconsistency , because the generation of security attack detectors is not always easily decomposable into a chain of thoughts, representing partial and incremental steps leading to the solution. Moreover, reasoning on the generated candidates may not be feasible, as the set of all the generated functions may not fit within the LLM’s context window. More importantly, our preliminary experiments suggest that LLMs struggle to assess the robustness of a set of candidate functions. Hence, we reuse only the general idea behind Self-Consistency, i.e., the LLM operating as a “self-ensemble”. Specifically, the non-determinism of LLMs is leveraged to produce a set of candidates that are assessed by the LLM itself, which generates an assessment dataset instead of directly assessing its own output. We refer to our novel variant of Self-Consistency as Self-Ranking: we ask the LLM to generate a synthetic dataset that simulates the presence of ground-truth data to select the best function. This synthetic dataset is used to evaluate and rank the generated functions.
@@
Once the synthetic dataset is generated, it serves as a testing ground for selecting the best function. Specifically, when multiple functions are generated, Self-Ranking starts evaluating each function with the synthetic dataset, ordering them based on a chosen metric, e.g., F2-Score, and retaining only the top-performing subset, denoted as the top_k functions. In our study, the usage of Self-Ranking aims to select the subset of $k$ most robust attack detectors.
@@ 4.1.2 Scenario TDA (Training Dataset Available)
In this scenario, a task-specific labeled training set of attacks is available to developers. This addition would allow them to directly train Machine Learning (ML) models that perform the task assigned to the generated detectors, without actually generating any function at all. In fact, SOTA techniques for many security tasks, including XSS and SQLi detection, use ML models trained on a labeled dataset.
@@ 4.2 Research Questions
  • RQ4. Transferability: Can the optimal parameters for function and synthetic data generation in one task be transferred and applied to achieve effective results in other tasks?
@@ 4.3.1 Models
We employ nine LLMs as listed in Table 1, which includes the aliases that are used throughout the remainder of this paper. Additionally, we report the Pass@1 scores from HumanEval chen2021evaluating as a proxy for their general reasoning capabilities: GPT-4T showed the best performance, followed closely by Opus, whereas PaLM exhibited the lowest performance among the models we considered. While Pass@1, which assesses the model’s reasoning capabilities to generate complex algorithms, may serve as an indicator of its expected performance on our task, we do not expect a perfect correlation with our results, since our task is more focused on utilizing extensive knowledge rather than reasoning on a given problem.
@@ 4.3.2 Configurable Parameters
[…]
Table 2 shows the domains of the configurations for the different parts of the empirical study. Given the domain, some of its configurations may fail on a given task. Hence, they are excluded. For example, among the configurations for Code Generation (see Table 2), when generating an XSS detector, all the configurations with models Sonnet, PaLM, Llama, Mixtral, and temperature 0 fail and are hence discarded. Similarly, for the SQLi detector, the aforementioned configurations, along with PaLM, Sonnet, and Mixtral at temperature 0.5, fail and are also excluded. Haiku and GPT-3.5T are completely discarded in code generation for the same reasons, but they are still among the employed models because they are used for synthetic dataset generation.
@@
Code Generation GPT-4T, GPT-4, Opus, Sonnet, PaLM, Llama, Mixtral 0.0, 0.5, 0.1 0, 2, 6, 10 True, False
Dataset Generation GPT-4T, GPT-3.5T, Opus, Sonnet, Haiku 0.0, 0.5, 0.1 0, 2, 6, 10 True, False
@@
For synthetic dataset generation, we iteratively generate a dataset of 100 examples, with an equal split of 50 malicious and 50 benign cases. We use a timeout of 9,000 seconds to discard configurations that fail to fill the dataset with 100 examples in a reasonable amount of time. To account for non-determinism, we repeat the process 10 times and produce 10 distinct datasets. The performance used to determine the ranking is calculated by averaging the performance on the 10 different produced datasets. When Self-Ranking is used, we have to decide the number $k$ of top-ranked functions to consider. In our experiments, the values of $k$ for the selection of top_k functions are 1, 3, and 5.
Several RAG sources were analyzed with preliminary experiments to select the most suitable ones for our setup. Considering only one source of knowledge for every task is reasonable since the focus of the empirical study is to exploit a RAG source under the assumption that it is available and of high quality, not to automatically find the best possible source of knowledge among the available ones. For the XSS detection task, we used the XSS Evasion Cheat Sheet by OWASP [^6] https://cheatsheetseries.owasp.org/cheatsheets/XSS_Filter_Evasion_Cheat_Sheet.html as the RAG source. OWASP is a foundation that works to improve the security of software and the knowledge contained in their cheatsheets is potentially useful to improve the quality of the detectors generated for XSS attack. For the SQLi detection task, the selected RAG source is an article on WebSec [^7] https://websec.wordpress.com/2010/12/04/sqli-filter-evasion-cheat-sheet-mysql/ containing a list of SQLi patterns. The knowledge shared in both sources is well-suited to our goal of creating effective detection functions via LLMs.
@@ 4.4.2 Dataset Generation
[…]

### v2

@@ Abstract
Large Language Models (LLMs) are increasingly used in software development to generate functions, such as attack detectors, that implement security requirements. Ensuring the LLMs have enough knowledge to address specific security requirements, such as information about existing attacks, is a key challenge. For this, we propose an approach integrating Retrieval Augmented Generation (RAG) and Self-Ranking into the LLM pipeline. RAG enhances the robustness of the output by incorporating external knowledge sources, while the Self-Ranking technique, inspired by the concept of Self-Consistency, generates multiple reasoning paths and creates ranks to select the most robust detector. Our extensive empirical study targets code generated by LLMs to detect two prevalent injection attacks in web security: Cross-Site Scripting (XSS) and SQL injection (SQLi). Results show a significant improvement in detection performance while employing RAG and Self-Ranking, with an increase of up to 71%pt [^1] Percentage points are the standard unit of measure for differences between percentages. (on average 37%pt) and up to 43%pt (on average 6%pt) in the F2-Score for XSS and SQLi detection, respectively.
@@ 1 Introduction
To this end, we adopt two components to the LLM pipeline: Retrieval Augmented Generation (RAG) lewis2020retrieval and Self-Ranking. RAG, a technique widely investigated in Natural Language Processing (NLP), enhances the quality of the output by incorporating external information. We use RAG to leverage existing knowledge sources that document known attacks to increase the robustness of the detectors generated by LLMs. Additionally, LLMs may exhibit non-deterministic behaviour in generation ouyang2023llm , which results in multiple, diverse solutions for the same prompt. We propose Self-Ranking, building upon the concept of Self-Consistency wang2023selfconsistency : we take advantage of the multiple reasoning paths associated with LLM’s non-determinism to rank the alternative outputs and select the best one. More concretely, the LLM can propose different implementations obtained querying the model multiple times, with Self-Ranking we propose a strategy to rank the different implementations and to keep only the best-performing ones. In turn, ranking is based on the creation of a synthetic dataset, also generated by LLMs, used to automatically assess the robustness of each candidate output.
@@
In our empirical study, we target two well-known and prevalent attacks: XSS and SQLi HYDARA2015170 ; lawal2016systematic , which are ranked second and third in ‘2023 CWE Top 25 Most Dangerous Software Weaknesses’ [^3] https://cwe.mitre.org/top25/archive/2023/2023_top25_list.html respectively. Our extensive empirical study involves the evaluation of nine different LLMs. Results indicate that integrating external knowledge with RAG improved detection performance up to 66%pt (on average 17%pt) on the F2-Score for XSS detection and up to 67%pt (on average 7%pt) for SQLi detection compared to relying solely on few-shot examples. F2-Score is a metric which gives double importance to recall than to precision, more details about this metrics and the reasons behind its choice will be discussed in Section 4.6. Additionally, employing Self-Ranking enhanced the LLM performance by up to 71%pt (on average 37%pt) on the F2-Score for XSS detection and up to 43%pt (on average 6%pt) for SQLi detection.
@@
State-of-the-art (SOTA) Machine Learning (ML) based attack detectors require a labeled training set of attacks, while our LLM-based approach is applicable even when no training set is available, provided a RAG source is accessible. In our empirical study, we considered also the scenario when a training set is available and SOTA methods can be used. In such a scenario, our approach achieved a performance comparable to SOTA XSS and SQLi machine learning-based detection methods, with the remarkable advantage that LLMs are pre-trained, while specialized machine learning-based methods require dedicated model design and training: developers can obtain SOTA performance by just querying an LLM using our RAG and Self-Ranking augmented pipeline. Another key advantage is that LLMs generate code that can be understood and manipulated by developers, while the decision making logic of a machine learning model is to a large extent opaque to developers. We also demonstrated the transferability capabilities of LLMs, which requires just configuration steps: after configuring the LLM (i.e., setting parameters such as the model checkpoint to use, the temperature, the number of few-shot examples) for optimal performance on one attack detection task (e.g., XSS), the resulting configuration can be transferred to another task (e.g., SQLi), with 16%pt and 10% improvement when transferring the LLM parameters from one task to another, as compared to average cases without parameter transfer. On the other hand, transferability between tasks for the SOTA detectors would require additional training steps, as well as a dataset specific to the target vulnerabilities.
@@
  • We introduce an approach that integrates RAG and Self-Ranking to evaluate and improve the robustness of LLM-generated attack detectors.
@@ 3.2 Few-shot Examples
>>>detect_xss(’http://anonymised-site/moderate-social-media&t=1396533765893&n=1129109&k=mainentity’)
@@
>>>detect_xss(’http://anonymised-site/guestbook/index.php?lang="><script>alert(document.cookie)</script>...’)
@@ 3.3 RAG
We followed the general RAG methodology, starting from the indexing phase. Once the knowledge sources about attacks are stored in vector representation [^5] The selection of appropriate data sources and their storage is discussed in Section 4 as it is an experimental choice., they can be queried to retrieve the most relevant information chunks (referred to as ‘External Knowledge’ in Figure 1). We utilize the Task (‘Detector Signature’ in Figure 1) directly as a query to retrieve these chunks, which are then appended to the Template to obtain a RAG Prompt, as illustrated in Listing 5. We chose to append the retrieved chunks to the template. While external knowledge could be added to the function’s comments, we found this approach to be less fluent and practical, especially given the size of the retrieved knowledge.
@@ 3.4 Self-Ranking
Self-Consistency does not apply directly to our task as proposed originally wang2023selfconsistency , because the generation of security attack detectors is not always easily decomposable into a chain of thoughts, representing partial and incremental steps leading to the solution. Moreover, reasoning on the generated candidates may not be feasible, as the set of all the generated functions may not fit within the LLM’s context window. More importantly, our preliminary experiments suggest that LLMs struggle to assess the robustness of a set of candidate functions. To explore this further, we applied the LLM-as-a-judge approach zheng2023judging , prompting GPT-4 to rate the robustness of several sampled detectors on a scale from 0 to 4. Over 95% of the scores fell within the top two categories (3 and 4), indicating poor discrimination among candidates. To assess whether these scores aligned with actual performance, we conducted a Spearman Correlation test between the LLM-assigned ranks and the F2-scores of the same detectors on a validation set. The correlation was negligible. A subsequent linear regression analysis, using the LLM score as the independent variable and F2-score as the dependent variable, yielded a 95% confidence interval ranging from 0.09-0.09 to $0.11$, reinforcing the conclusion that no meaningful relationship exists. These findings highlight the inadequacy of direct LLM-based evaluation for robustness scoring. Hence, we reuse only the general idea behind Self-Consistency, i.e., the LLM operating as a “self-ensemble”. Specifically, the non-determinism of LLMs is leveraged to produce a set of candidates that are assessed by the LLM itself, which generates an assessment dataset instead of directly assessing its own output. We refer to our novel variant of Self-Consistency as Self-Ranking: we ask the LLM to generate a synthetic dataset that simulates the presence of ground-truth data to select the best function. This synthetic dataset is used to evaluate and rank the generated functions.
@@
Once the synthetic dataset is generated, it serves as a testing ground for selecting the best function. Specifically, when multiple functions are generated, Self-Ranking starts evaluating each function with the synthetic dataset, ordering them based on a chosen metric, e.g., F2-Score, and retaining only the top-performing subset, denoted as the top_k functions. In our study, the usage of Self-Ranking aims to select the subset of $k$ most robust attack detectors. Unlike the LLM-as-a-judge methodology, our Self-Ranking approach does not ask one LLM to directly judge another’s output. Instead, we use synthetic data generated by an LLM to evaluate the performance of our detectors.
@@ 4.1.2 Scenario TDA (Training Dataset Available)
In this scenario, a task-specific labeled training set of attacks is available to developers. This addition would allow them to directly train ML models that perform the task assigned to the generated detectors, without actually generating any function at all. In fact, SOTA techniques for many security tasks, including XSS and SQLi detection, use ML models trained on a labeled dataset.
@@ 4.2 Research Questions
  • RQ4. Transferability: Can the optimal parameters for function and synthetic data generation in one task be transferred and applied to achieve effective results in the other task?
@@ 4.3.1 Models
We employ nine LLMs as listed in Table 1, which includes the aliases that are used throughout the remainder of this paper. At the time we started the experiments, we selected the best performing models available to us through the SAP AI Core service, which allowed us access a broad range of state of the art LLMs. We further filtered the available models based on their performance on the HumanEval chen2021evaluating benchmark. We also consider some of the weaker models to better highlight the differences of results in the considered task. Additionally, we report the Pass@1 scores from HumanEval as a proxy for their general reasoning capabilities: GPT-4T showed the best performance, followed closely by Opus, whereas PaLM exhibited the lowest performance among the models we considered. While Pass@1, which assesses the model’s reasoning capabilities to generate complex algorithms, may serve as an indicator of its expected performance on our task, we do not expect a perfect correlation with our results, since our task is more focused on utilizing extensive knowledge rather than reasoning on a given problem.
@@ 4.3.2 Configurable Parameters
A model configuration refers to the model and temperature used (e.g., GPT-4 with temperatures $0.5$ and $1.0$). A prompt configuration is determined by the number of Few-shot examples (with equal malicious and benign examples) and the choice to use RAG or not, to enrich the prompt. By combining model and prompt configurations, we obtain an overall Configuration Conf, consisting of a tuple with a model, a temperature, the few-shot number, and the RAG usage choice. For instance, $(\textsc{GPT-4},0.5,2,T)$ indicates the configuration where the model is GPT-4, the temperature is set to 0.5, the number of Few-shot examples is 2, and RAG is used (T represents True for the RAG Usage Parameter).
Table 2 shows the domains of the configurations for the different parts of the empirical study. Given the domain, some of its configurations may fail on a given task. Hence, they are excluded. By failure, we mean that the generation process did not produce a valid artifact. Specifically, for Code Generation, failure occurs when more than 80% of the generated functions raise exceptions upon execution. For Dataset Generation, a configuration fails if it cannot produce a valid CSV file meeting the specified requirements (e.g., number of samples and class distribution) within three hours. For example, among the configurations for Code Generation (see Table 2), when generating an XSS detector, all the configurations with models Sonnet, PaLM, Llama, Mixtral, and temperature 0 fail and are hence discarded. Similarly, for the SQLi detector, the aforementioned configurations, along with PaLM, Sonnet, and Mixtral at temperature 0.5, fail and are also excluded. Haiku and GPT-3.5T are completely discarded in code generation for the same reasons, but they are still among the employed models because they are used for synthetic dataset generation.
@@
Code Generation
GPT-4T, GPT-4, Opus,
Sonnet, PaLM,
Llama, Mixtral
0.0, 0.5, 0.1 0, 2, 6, 10 True, False
Dataset Generation
GPT-4T, GPT-3.5T, Opus,
Sonnet, Haiku
0.0, 0.5, 0.1 0, 2, 6, 10 True, False
@@
For synthetic dataset generation, we iteratively generate a dataset of 100 examples, with an equal split of 50 malicious and 50 benign cases. Specifically, we prompt the LLMs to generate datasets in this format, although we do not enforce any hard constraints to guarantee strict adherence to the desired class distribution. Because of that, we do not have any guarantee about the correctness of the labels of the samples produced by the LLMs. We use a timeout of 9,000 seconds to discard configurations that fail to fill the dataset with 100 examples in a reasonable amount of time. To account for non-determinism, we repeat the process 10 times and produce 10 distinct datasets. The performance used to determine the ranking is calculated by averaging the performance on the 10 different produced datasets. When Self-Ranking is used, we have to decide the number $k$ of top-ranked functions to consider. In our experiments, the values of $k$ for the selection of top_k functions are 1, 3, and 5.
Several RAG sources were analyzed with preliminary experiments to select the most suitable ones for our setup. Considering only one source of knowledge for every task is reasonable since the focus of the empirical study is to exploit a RAG source under the assumption that it is available and of high quality, not to automatically find the best possible source of knowledge among the available ones. For the XSS detection task, we used the XSS Evasion Cheat Sheet by OWASP [^6] https://cheatsheetseries.owasp.org/cheatsheets/XSS_Filter_Evasion_Cheat_Sheet.html as the RAG source. OWASP is a foundation that works to improve the security of software and the knowledge contained in their cheatsheets is potentially useful to improve the quality of the detectors generated for XSS attack. For the SQLi detection task, the selected RAG source is an article on WebSec [^7] https://websec.wordpress.com/2010/12/04/sqli-filter-evasion-cheat-sheet-mysql/ containing a list of SQLi patterns. The knowledge shared in both sources is well-suited to our goal of creating effective detection functions via LLMs. To further elaborate on the preliminary experiments conducted to select the RAG sources, we manually curated a set of sources that we considered reliable. In addition to the ones ultimately selected, for XSS detection we considered the OWASP Web Security Testing Guide [^8] https://owasp.org/www-project-web-security-testing-guide/latest/4-Web_Application_Security_Testing/07-Input_Validation_Testing/01-Testing_for_Reflected_Cross_Site_Scripting and a widely-read Invicti article on XSS filter evasion [^9] https://www.invicti.com/blog/web-security/xss-filter-evasion/. For SQL injection detection, we additionally evaluated the OWASP SQL Injection Prevention Cheat Sheet [^10] https://cheatsheetseries.owasp.org/cheatsheets/SQL_Injection_Prevention_Cheat_Sheet.html and the W3Schools page on SQL injection [^11] https://www.w3schools.com/sql/sql_injection.asp. We opted not to include these alternative sources in the main experimental configuration to avoid a combinatorial explosion in the number of experiments. Instead, we evaluated their performance by measuring the F2 scores on a representative configuration sample $(\textsc{GPT-4},0.5,2,T)$, selecting the best-performing source on the val_setval\_set. For XSS detection, the two alternatives performed 6%pt and 24% worse than the selected source. Similarly, for SQLi detection, the OWASP Cheat Sheet and W3Schools performed 25%pt and 33% worse, respectively, compared to the chosen source.
@@ 4.4.1 Function Generation
We consider the generation of duplicates to be natural when LLM deal with code generation. Specifically, duplicates are more frequent at lower temperatures. We chose not to remove duplicates because they may reflect the LLM’s consistency and confidence in its output. In this context, repeated solutions can be interpreted as an indication that the model assigns higher confidence to those particular outputs. This approach is consistent with established evaluation metrics such as pass@k chen2021evaluatinglargelanguagemodels , which assesses the functional correctness of generated code by considering diverse solutions but does not include any mechanism to handle or account for duplicate outputs.
@@ 4.4.2 Dataset Generation
To account for non-determinism, we repeat the generation process $m$ (in our experiments, 10) times and obtain a Synthetic Dataset Run $S(t,C)=\{s_{1}(t,C),\ldots,s_{m}(t,C)\}$. In Figure 2, $m=3$ datasets are generated, $s_{1},s_{2},s_{3}$ for each configuration. Given all the configurations in the domain, we define a Synthetic Dataset Experiment $P(t)$ as $\{S(t,Conf_{1})$, $S(t,Conf_{2})$, $\ldots\}$. In Figure 2, the experiment $P$ contains two sets of generated datasets, $S_{1}$ and $S_{2}$.